# From Asymptotically Flat Gravity to Finite Causal Diamonds

### v1

@@ Abstract
We demonstrate that the phase space of the soft sector of asymptotically flat gravity in four spacetime dimensions can be identified with that of a spherically symmetric finite casual diamond in Minkowski spacetime. The leading soft graviton mode is identified with the radial fluctuation of the causal diamond size, while the Goldstone mode involves both the radial fluctuation and its symplectic partner. This allows us to relate the radial fluctuations of the causal diamond with the asymptotic transverse fluctuations parametrized by the soft modes.
@@
$\begin{split}\Delta a_{{\text{\tiny AFG}}}(u,z,{\bar{z}})&\equiv a_{{\text{\tiny AFG}}}(u,z,{\bar{z}})-\pi(u_{0}-u)^{2}\\ &=-8\pi\int_{u}^{\infty}{\text{d}}u\,m_{B}(u^{\prime},z,{\bar{z}}).\end{split}$ (30)
@@
$\begin{split}\Delta A_{{\text{\tiny AFG}}}(u)&\equiv\int_{S^{2}}{\text{d}}^{2}z\,{\gamma}_{z{\bar{z}}}\Delta a_{{\text{\tiny AFG}}}(u,z,{\bar{z}})\\ &=-32\pi^{2}\int_{u}^{\infty}{\text{d}}u\,\bar{m}_{B}(u^{\prime}),\end{split}$ (31)
@@
${\epsilon}=-16\pi\bar{m}_{B}(-\infty).$ (34)
@@
$\begin{split}m_{B}(u,z,{\bar{z}})&=\frac{1}{8}\Box(\Box+2)N(z,{\bar{z}})\big(1-\theta(u-u_{s})\big).\end{split}$ (37)
@@
$\begin{split}\bar{m}_{B}(-\infty)=\frac{1}{2}\bar{N},\end{split}$ (38)
@@
$\begin{split}\bar{N}=-\frac{1}{8\pi}{\epsilon}\,.\end{split}$ (39)
@@
$\begin{split}\bar{C}=-16\pi\mu L+h(L),\end{split}$ (41)
@@
$\begin{split}\bar{P}_{-}&\equiv\frac{1}{4\pi}\int_{S^{2}}{\text{d}}^{2}z\,{\gamma}_{z{\bar{z}}}P_{-}=\frac{1}{8\pi G}\bar{N}=-\frac{{\epsilon}}{G}.\end{split}$ (46)
@@
Acknowledgments. We would like to thank Mathew Bub, Sucheta Majumdar, Prahar Mitra, and Sabrina Pasterski for useful discussions. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. L.C. is supported by the Celestial Holography Simons collaboration. T.H. and K.Z. are supported by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2817, the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award No. DE-SC0011632, and the Walter Burke Institute for Theoretical Physics. K.Z. is also supported by a Simons Investigator award.

### v2

@@ Abstract
We demonstrate that the phase space of the soft sector of asymptotically flat gravity in four spacetime dimensions can be identified with that of a spherically symmetric finite casual diamond in Minkowski spacetime. The leading soft graviton mode is geometrically identified with the radial fluctuation of the causal diamond size, while the Goldstone mode involves both the radial fluctuation and its symplectic partner. This allows us to relate the radial fluctuations of the causal diamond with the asymptotic transverse fluctuations parametrized by the soft modes.
@@
$\begin{split}\Delta a_{{\text{\tiny AFG}}}(u,z,{\bar{z}})&\equiv a_{{\text{\tiny AFG}}}(u,z,{\bar{z}})-\pi(u_{0}-u)^{2}\\ &=-8\pi\int_{u}^{\infty}{\text{d}}u^{\prime}\,m_{B}(u^{\prime},z,{\bar{z}}).\end{split}$ (30)
@@
$\begin{split}\Delta A_{{\text{\tiny AFG}}}(u)&\equiv\int_{S^{2}}{\text{d}}^{2}z\,{\gamma}_{z{\bar{z}}}\Delta a_{{\text{\tiny AFG}}}(u,z,{\bar{z}})\\ &=-32\pi^{2}\int_{u}^{\infty}{\text{d}}u^{\prime}\,\bar{m}_{B}(u^{\prime}),\end{split}$ (31)
@@
${\epsilon}=-16\pi\bar{m}_{B}(-\infty),$ (34)
where the equality is implicitly understood to mean that the two sides are equal under an isomorphism between the two phase spaces which preserves the brackets.
@@
$\begin{split}m_{B}(u,z,{\bar{z}})&=-\frac{1}{8}\Box(\Box+2)N(z,{\bar{z}})\big(1-\theta(u-u_{s})\big).\end{split}$ (37)
@@
$\begin{split}\bar{m}_{B}(-\infty)=-\frac{1}{2}\bar{N},\end{split}$ (38)
@@
$\begin{split}\bar{N}=\frac{1}{8\pi}{\epsilon}\,.\end{split}$ (39)
@@
$\begin{split}\bar{C}=16\pi\mu L+h(L),\end{split}$ (41)
@@
$\begin{split}\bar{P}_{-}&\equiv\frac{1}{4\pi}\int_{S^{2}}{\text{d}}^{2}z\,{\gamma}_{z{\bar{z}}}P_{-}=\frac{1}{8\pi G}\bar{N}=\frac{{\epsilon}}{G}.\end{split}$ (46)
@@
Acknowledgments. We would like to thank Mathew Bub, Marc Klinger, Sucheta Majumdar, Prahar Mitra, and Sabrina Pasterski for useful discussions. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities. L.C. is supported by the Celestial Holography Simons collaboration. T.H. and K.Z. are supported by the Heising-Simons Foundation “Observational Signatures of Quantum Gravity” collaboration grant 2021-2817, the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award No. DE-SC0011632, and the Walter Burke Institute for Theoretical Physics. K.Z. is also supported by a Simons Investigator award.